# A Methodological Report on Anomaly Detection on Dynamic Knowledge Graphs

### v2

@@ 3.3 Two-Hop Graph Structure
Another method of capturing two-hop neighbours is using the Self-Attention mechanism (SA) to capture the correlations among vertices. SA works by computing the relevance (or attention) between each pair of elements in a sequence, enabling the model to weigh the importance of each element relative to others. This allows the model to focus on relevant parts of the input sequence when making predictions, effectively capturing long-range dependencies and complex relationships within the data. See the "two hop graph structure" part in Fig. 2.
@@ 3.4 Ensemble Learning
Ensemble learning is a prevalent methodology in machine learning that reduce variance and enhance model robustness by combining multiple models. The effectiveness of ensemble learning depends on the integration of diverse, high-performing models [14]. It allows multiple models to make predictions on the same problem and vote on the final predictions. Voting can be hard or soft voting, with the mechanism such as unanimous and majority voting, which will be introduced in this subsection.
Figure 3: The ensemble learning framework used in this competition. The best-performing models: XGBoost (XGB), Support Vector Machine (SVM), and Self-Attention (SA) are combined using three voting mechanisms: hard-unanimous, hard-majority, and soft voting, to generate the prediction $\hat{y_{1}}$, where $\hat{y_{1}}\in\{0,1\}$, with 0 indicating normal and 1 indicating anomaly. Optionally, a two-stage model combination can be applied, where the unsupervised Isolation Forest further refines the anomalies detected in $\hat{y_{1}}$ to reduce false positives.
@@ 4.2 Performance of Models
This section presents the experimental results for the three data representations: D1, D2, and D3, correspond respectively to sequential data, one-hop, and two-hop graph representations. Table 2 summarises the performance of individual models, while table 3 focuses on the results of different ensemble methods.
The first part of the experiments evaluates the performance of each model under different data representations, as shown in Table 2. The results indicate that incorporating graph structures and service interactions generally enhances model performance, particularly in recall and F1 score.
@@
The second part of the experiments evaluates the performance of different ensemble methods, as shown in table 3. We tested combinations of XGB, SVM, Self-Attention, and Isolation Forest, along with different voting mechanisms: soft, hard-unanimous, and hard-majority. Table 3 shows that the best performance is achieved with the combination of SVM, XGB and SA in hard-unanimous voting.
@@
Figure 4: The figure shows the F1 scores of the models across the three dataset levels, as well as the F1 scores obtained after applying ensemble learning methods. The individual models include XGB, SVM, TCN, LSTM, GRU, and the Self-Attention Model. The models used in the ensemble learning approach are XGB, Isolation Forest, SVM, and the Self-Attention Model.
Fig. 4 shows that the one-hop representation, two-hop representation as well as the ensemble learning methods all significantly outperforms the baseline result  [^5] The baseline result is from https://eval.ai/web/challenges/challenge-page/2267/overview, demonstrating the potential of our approach.
@@ 5 Conclusion
In this paper, we explored the effectiveness of various machine learning and deep learning models for anomaly detection on dynamic knowledge graphs, focusing on a microservices environment. Using a structured experimental approach, we analysed the performance of the models on three representations: sequential data, one-hop, and two-hop graph structures.
Our results show that incorporating graph structures significantly enhances model performance. The two-hop graph structure, which includes service interactions, delivered the best results across most metrics, emphasising the importance of capturing complex interdependencies. Deep learning models, especially the Self-Attention model, performed particularly well in leveraging these structures.
Furthermore, the ensemble methods significantly boosted performance by combining the strengths of the top individual models. Unanimous voting, which requires all models to agree on an anomaly, outperformed majority voting in terms of precision, matching the competition’s emphasis on precision.
Future work should focus on several improvements. First, improving model scalability to handle larger, more complex datasets is essential. Second, exploring advanced ensemble techniques such as stacking or blending could improve detection accuracy. Third, integrating real-time anomaly detection would benefit practical applications in dynamic environments. Additionally, our study was limited to binary classification; future research should improve the accuracy of detecting different types of anomalies. Finally, although we used graph structure information, the application of graph neural networks remains a promising direction for further investigation.

### v3

@@ 3.4 Hierarchical and Inter-Service Representation
Another method is adding implicitly the inter-service dependencies using the model layers, like Self-Attention, LSTM to capture the correlations among vertices during the model training. SA works by computing the relevance (or attention) between each pair of elements in a sequence, enabling the model to weigh the importance of each element relative to others. This allows the model to focus on relevant parts of the input sequence when making predictions, effectively capturing long-range dependencies and complex relationships within the data. See the "Inter-Service Dependencies" part in Fig. 4.
@@ 3.5 Ensemble Learning
Ensemble learning is a prevalent methodology in machine learning that reduce variance and enhance model robustness by combining multiple models. The effectiveness of ensemble learning depends on the integration of diverse, high-performing models [14]. It allows multiple models to make predictions on the same problem and vote on the final predictions. Voting can be hard or soft voting, with mechanisms such as unanimous and majority voting, which will be introduced in this subsection.
Figure 5: The ensemble learning framework used in this competition. The best-performing models: XGB, SVM, and SA are combined using three voting mechanisms: hard-unanimous, hard-majority, and soft voting, to generate the prediction $\hat{y_{1}}$, where $\hat{y_{1}}\in\{0,1\}$, with 0 indicating normal and 1 indicating anomaly. Optionally, a two-stage model combination can be applied, where the unsupervised Isolation Forest further refines the anomalies detected in $\hat{y_{1}}$ to reduce false positives.
@@
In the pipeline, we also explore an optional two-stage model to improve anomaly detection accuracy. In this approach, we first identify potential anomalies by merging the results of the three models through hard or soft voting, and then use these initial predicted anomalies as a new dataset $\mathbf{X}_{2}$, and subsequently feed $\mathbf{X}_{2}$ into the unsupervised algorithm Isolation Forest as a secondary filtering mechanism.
The aim of incorporating IF into this two-stage pipeline is to reduce the number of false positives through tree-based segmentation and isolation of outliers. This approach allows us to identify unique, isolated anomalies that may have been inaccurately classified in the first stage, thereby improving the robustness of the overall anomaly detection process.
@@ 4.2 Performance of Models
This section presents the experimental results for the three data representations: D1, D2, and D3, correspond respectively to sequential data, hierarchical data, and inter-service data. Table 2 summarises the performance of individual models, while table 3 focuses on the results of different ensemble methods.
The first part of the experiments evaluates the performance of each model under different data representations, as shown in Table 2. Table 2 only showcases the best results for each case. It is worth remarking that the baseline score provided by the challenge organisers is the score obtained by predicting all snapshots given at each timestamp as anomalies, and thus it has a recall equal to 1. The results indicate that incorporating graph structures and service interactions generally enhances model performance, particularly in recall and F1 scores.
While our study aimed to evaluate multiple models comprehensively across different configurations, certain constraints affected the depth of testing. Due to daily submission limits, some models could not be exhaustively tested across all configurations. For instance, XGB and SVM were only evaluated under configuration D2. D3 representation needs to incorporate additional inter-service information, which requires significant computational resources, limiting our ability to test them further. Moreover, the TCN model detected no anomalies following training under D3; as a result, its findings are omitted from this analysis.
@@
The second part of the experiments evaluates the performance of different ensemble methods, as shown in table 3. Table 3 only shows the best results for each case. We tested combinations of XGB, SVM, SA, and IF, along with different voting mechanisms: soft, hard-unanimous, and hard-majority. The results show that the combination of SVM, XGB and SA achieves the best performance in hard-unanimous voting.
@@
Figure 6: The figure shows the F1 scores of the models across the three data representations in the left, as well as the F1 scores obtained after applying ensemble learning methods in the right. The individual models include XGB, SVM, TCN, LSTM, GRU, and SA Model. The models used in the ensemble learning approach are XGB, IF, SVM, and SA Model.
Fig. 6 shows that the hierarchical data representation, inter-service dependency representation as well as ensemble learning methods all significantly outperforms the baseline result, demonstrating the potential of our approach.
@@ 5 Conclusion
In this paper, we explored the effectiveness of various machine learning and deep learning models for anomaly detection on dynamic knowledge graphs, focusing on a Micro-services environment. Using a structured experimental approach, we analysed the performance of the models on three representations: sequential data, hierarchical data, and inter-service dependency data.
Our results show that incorporating graph structures significantly enhances model performance. The inter-service structure, which includes service interactions and hierarchical representation, delivered the best results across most metrics, emphasising the importance of capturing complex structures and (inter-)dependencies. Deep learning models, especially the SA model, performed particularly well in leveraging these structures.
Furthermore, the ensemble methods significantly improved performance by combining the strengths of the top individual models. Unanimous voting, which requires all models to agree on an anomaly, outperformed majority voting in terms of precision, matching the competition’s emphasis on precision.
Future work should focus on several improvements. First, improving model scalability to handle larger, more complex datasets is essential. Second, exploring advanced ensemble techniques such as stacking could improve detection accuracy. Third, integrating real-time anomaly detection would benefit practical applications in dynamic environments. Additionally, our study was limited to binary classification; future research should improve the accuracy of detecting different types of anomalies. Finally, although we used graph structure information, the application of graph neural networks remains a promising direction for further investigation. Moreover, our data representations are high-dimensional, we can compare and use some low-dimensional representations, such as node embedding, graph embedding to test our proposed solutions.